### v18

1. Introduction
The ideas here are meant to give a general procedure for calculating an optimal sufficient value for the number of labels required in a standard labeling problem. The techniques apply to vertex labeling and edge labeling problems. Usually, the calculation will give a rough estimate of the optimal sufficient value, rather than an exact value. However, the calculation should bound the optimal value from above in most applications.
There are several steps to proving the main result, which is that given three sufficient conditions, it is possible to calculate upper bounds on the number of labels required in the general labeling problem. The first sufficient condition is that there is a uniform map that carries the fibers to one another transitively. This condition requires that the order on the permutations of the fibers is In most cases, this last equality will follow from the phrasing of the problem, but in the case that the labeling is on the vertices of the graph, and duplications are counted on induced labels on the edges, and the fibers of induction are the edges of the graph under the labeling scheme, the equality for requires that the graph under the labeling scheme is acyclic. Then, in addition, there is a basis step to an inductive argument. Usually, this step is verified by applying Markovβs Inequality to an expectation that is calculated for each of the individual fibers, before performing the identification allowed by the existence of the uniform map. Once the basis step is verified, apply induction.
The induction implies that there is a simple test for the verification of a fixed upper bound on number of required labels: The required step here is verification of the reference set hypothesis; show that the reciprocal of the sample set for the number of labelings is less than times the probability that there is a labeling without collisions. The probability, based on an expectation parameterized by the projected required strength of the labeling, is calculated using integration given by the induction from the basis step of the induction. There is a correlation between the size of an expected value, and the existence of a labeling without collisions, in the framework of each individual labeling problem.
Look more closely at packings with exactly two elements, which are called -body problems. For most -body problems, depending on the phrasing of the problem, formulate results without using the required hypothesis that there is a placement of the two graphs so that the sum of the degree sequences is majorized by the degree sequence of the complete graph. There are other hypotheses used as substitutes for the degree packing hypothesis. In these cases, there is a -body allowance that permits the use of without the natural doubling of expectation that is normally required in figuring
(Example. The Bollobas-Eldridge Conjecture) The Bollobas-Eldridge Conjecture is formally stated that if then and have an edge-disjoint placement in . There is a very strong analogue. There are complications in every step of the method with regards to analyzing the proof of the result. All the complications and some of the nuances of the method in this setting are explained next. The maps have to allow label replacement to get the result. So, there is a penalty for vertex duplication in the expectation. Step 3 is complicated by the assessment of the penalty for vertex label duplication and in the check of the reference set hypothesis. Additionally, the calculation for the value of is somewhat complicated.
There are some obvious extremal examples such as Furthermore, the examples of disjoint collections of suitably ordered complete graphs are extremal in terms of the problem statement. There is a previous outline of an approximation of the result with previous ideas: the reference set method shows an approximation of the Sauer-Spencer criteria. The previous result can be improved to show a closer approximation of the desired conjecture in most cases. Outline the problem as a vertex labeling problem (as opposed to an edge labeling problem where the edge label duplications are counted at the vertex fibers) and form where the terms are the expected number of edge label duplications between and in the problem. Additionally, in this case, is set
Step 1. Use a map from to the looped complete graph where and a map from to the same target. The symmetric map from into is uniform, as is the map from The map allows replacement; the application requires that the maps be injections, and therefore, there is a penalty assessed on the random variable and the expectation of the random variable. Checking the degree sequence requirement is not required in this case, because there are only two graphs in the packing.
Step 2. There are two parts to the stage in the problem. There are two parts to . The first part comes from the expectation in the problem. The second part is a penalty for the duplication of vertex labels. Consider all the placements of and in where replacement is allowed on the vertex labels of and (In this way the problem is highly atypical because loops are allowed and then counted in the random variable.) In this case is the expected number of edge label intersections in closed neighborhoods from against closed neighborhoods of the vertices taken over the range of and Count the number of ways edges by dividing into three parts and considering all the separate cases from various sets of closed vertex neighborhoods: Consider first the vertices are mapped to the same target in and Consider and Take This should be clear from simple counting. Use in the denominator of the multiplier instead of because the labeling allows replacement, and so the product is uniform over choices instead of . The degree of the vertices is instead of as in a standard complete graph. There is a penalty for vertex label duplication, but it is not applied in this part of the expectation. Consider second the total expectation that both and Compare the closed neighborhood of and This expectation is The counting here is more complicated, but most of the main ideas can be gathered from the expression for . Consider third the vertices are mapped to the same target in and Consider and The third component of the expectation is The counts appear correct.
Set Sum over all ordered pairs of vertices from and respectively, and each vertex labeling is counted twice, once for each endpoint of the edge. Notice the allowance for the -body problem. Then there is a penalty for allowing vertex-replacement: Take the penalty equal to Do not double the penalty for the two elements of every duplication, and there is no natural doubling because the fibers are vertices. The -body allowance does not apply to the penalty. The number of fibers and the comparison orbit of the fibers in the penalty have identical cardinality. Take Count the vertices of as fibers, and each fiber has the comparison orbit cardinality (Do not count the vertex fibers as ) These are most of the difficult steps.
Step 3. In the reformulation of the problem, this step holds with and Take and use Markovβs Inequality. See that if then the basis step is satisfied. To check the reference set hypothesis, use . Count the number of labelings with replacement of both graphs. Then is about and is so this checks asymptotically. (It is required that be chosen slightly smaller than ) Try to use the edges as the fibers of induction, and get a function that is hyper-exponential in some base, and the form of is only
Continue by addressing two of the labeling problems that involve finding resolutions of trees in graphs and digraphs. The applications that follow appear mostly limited to trees.
(Example. Erdos-Sos) Here, the -body specification applies to the first part of the expectation. Both elements of the duplications on the vertex labels are internal to the first graph, so there is no -body allowance for the penalty portion of the expectation.
Step 1. The fibers of the problem are the edges. The labeling is on the vertices. The problem requires that the graph under resolution is a tree or forest. Because there are only two graphs in the packing, it is not necessary to verify the degree sequence condition. Use the -body allowance, and take the edges as the fibers of induction. It appears necessary to take a penalty in the expectation for vertex label duplication, and use
Step 2. Here, let the tree have order . Assume the host graph has average degree taken slightly larger than the requirement of the problem. Then Then for this component is and For the penalty portion, Then for this portion is applied as and the -body short-cut does not apply: The penalty is The penalty portion of the expectation doubles. So then
Step 3. There is a sufficiently large average degree in the host graph where the basis step will check. See that it is in the neighborhood of the specifications of the problem. Estimate the value of as using the expression for The idea is to show that Erdos-Sos is asymptotically accurate. The check of the reference set hypothesis is because in most cases is sufficiently smaller than . If beta is not sufficiently small, adjust the density of the graph slightly. Do not choose much smaller than because it is not required in the basis step of induction. Then the basis step and the reference set hypothesis both check, and these calculations give a rough form of the Erdos-Sos Conjecture.
(Example. Sumnerβs Conjecture) Here, the -body specification applies to the first part of the expectation, but not in the second part, because all the collisions on labels of the vertex set are internal to the problem.
Step 1. Skip the details for this step. The problem is phrased as a packing problem.
Step 2. Try to pack in the tournament Take Take Then Take, for the penalty Take The -body allowance does not apply to and so The penalty portion does not naturally double, but because the fibers have two endpoints, the penalty portion doubles.
Step 3. The basis step does not check unless the size of the tree is reset to and is sufficiently smaller than In these cases, the basis step checks because To check the reference set hypothesis, use the reference set because replacement is allowed.
For the -body problems, use the edges of the smaller graph as the fibers of induction; it is usually necessary to use the comparison orbit and maintain that the smaller graph in the placement is acyclic. In order to perform the basis step with use a comparison orbit of cardinality with an acyclic graph, to reconcile the basis step with an edge by edge resolution of a graph whose edges form the only fibers in the induction. Then, in all these cases, it is usually necessary to take the penalty for vertex label duplication. Try to extend the Erdos-Sos result to a -factor, using the edges as fibers, then the reference set hypothesis will not check unless The necessary restriction on in this example excludes any new results, and gives standard results within accepted parameters. Similarly, use the vertices as fibers and get results for cycles, and show gives and gives Try to extend the Sumner Conjecture to cycles. The phrasing of the problem as a vertex fiber problem will not permit a check of the reference set hypothesis.
(Example. Turanβs Theorem) Try deriving an approximation to Turanβs Theorem: The graph if and The closest approximation here is that packs in when the complement of has with In this case, do not apply the -body allowance; do not take the full penalty for vertex label duplication.
Step 1. This step is carried out using the symmetric map from the vertex set of to This map is uniform on the edges of the selected. Because the selected graph and the target graph are isomorphic to , it is clear the map will be uniform. This example is unusual because Do not check the degree sequence requirement in this example. Use the vertices as the fibers.
Step 2. The calculation of does not use the -body allowance. The constant of proportionality is 4 in this case. The duplication units are edges and the fibers are vertices. Use as the density of null edges in the graph that the vertices are chosen from. Use the formula to check the base case. (Pick about at a vertex of if is a small fraction close to and using exponentiation. This calculation appears to hold in general for cliques, but is not applied in this setting.) Take Then and and the value when See the case where the density of edges in the maximal complementary graph should be on the order of -factor: the choice of here as gives that is about
Step 3. Verify the basis step. Take the fibers of induction to be the vertices of Then take Again, choose and sufficiently small. Then is the calculation for the reference set hypothesis. By choice of sufficiently small choice is taken sufficiently smaller than
where and The calculation measures that the probability of a labeled graph in vertex labels intersected with and having no empty edges as greater than the reciprocal of the number of vertex labelings of with labels (where only vertices of are labeled). This would verify an approximate of Turanβs Theorem by the reference set hypothesis for suitably small The point is to drive smaller than because it is necessary to check the reference set hypothesis. Here and larger choices of give larger choices of The resultant larger choices of are more likely to disobey the reference set hypothesis. This makes intuitive sense because more null edges would decrease the likelihood of giving a choice of a complete The value is approximately the largest value that checks in the basis step. The value is in the denominator of the value for and is taken It has to be that is satisfied. The formulation of the theorem does not give information about the unique extremal graphs generated by the problem. Apply Markovβs Inequality in this context. Here, the choice of is limited by the basis case and the value of does not appear optimal in this problem.
Summarize some of the techniques that are available for edge-disjoint placement problems. First, to pack for generally check the degree sequence condition, use a covering space with a multigraph, and possibly use a transformation on the random variable. (In this case, the graphs have to be acyclic.) Then, to pack and there are several different formulations of the problem, most of which rely on the -body allowance in some way, and some of which do not require the degree sequence condition as an initial hypothesis. First, take the edges as the fibers of induction. In this case, take the penalty for vertex label duplication and use the comparison orbit Do not check the degree sequence requirement. To apply this case, the graph has to be acyclic. Next, use the idea from the example on the Bollobas-Eldridge Conjecture.
In this case, check the degree sequence requirement explicity, but add to the expectation to account for the vertex label duplication because of the phrasing of the problem. Third, use the vertices as fibers, with the added penalty for vertex label duplication, with similar conditions to the edge fiber case, so that neither graph has to be acyclic, but formulate in terms of and take Additionally, in this case, the penalty for vertex duplication is based on instead of the order of the graph Here, it is not necessary to check the degree sequence condition separately. Fourth, take the vertices of one graph as the fibers of induction, and skip checking the degree sequence requirement. In this case, set the comparison orbit equal to the product of the cardinalities of the vertex sets of the individual graphs:
This formulation does not appear to have any optimal applications. This last formulation also appears to apply in the case but again, does not seem to have any optimal applications.
2. Degree sequences and edge-disjoint placements
First, notice the following: Suppose the inequality is the only inequality available in proving the base case of Step 3. Then it may be that the smallest inequality that makes sense is actually for The strongest conclusion on using this basis step on the reference set method is
This is important, because it generally means problems cannot be solved with the sample space out of factorials by extending the support from to
There are additional complications when the problem is an edge-disjoint packing problem. There are alternative phrasings for edge-disjoint packing problems: One alternative is to use the fibers as edges and to use the entire range of the complete graph as the comparison orbit. Then and let where and The expectation
so that in the case that the packing is dense. Try the case of two spanning stars. Since the degree sequences do not pack, it is not possible to pack the two stars in an edge-disjoint fashion; try to verify that there is a packing in some of the other situations without checking the degree sequence condition. Use the full range of the edge set of the complete graph as the comparison orbit, or formulate the problem alternatively. One alternative is to use the vertices as fibers and to use the cardinality of the cross product of the vertex sets as the comparison orbit. Generally, there is the identical difficulty when the number of elements in the packing is in the case of two spanning stars, show that this parameterization will not give false positive results. Then and where and where the packing has elements. In the case of two stars, take and Then and clearly the method is impractical. There are similar difficulties when the problem is rephrased for different elements in the graph-packing. Both alternatives for the formulation of the random variable present obstacles that are realized to some extent by extremal examples. Consider the case of the star and the matching: In that case, given this parameterization, so that when is the order of the star. There are various ways of considering the same problem for packing sets of graphs. One of the goals is to set up the problem so that the ideas can be applied in a wide range of settings. The following idea for setting up the problem is more flexible, but has some added requirements that are difficult to satisfy.
The graph packing problem can be resolved in some cases by adding a degree sequence packing requirement. Proceed by taking and then labeling the resultant packing as a multigraph that is labeled and re-labeled. Let be some packing of a set of acyclic graphs Generally work in instead of to avoid problems. Take as usual. Suppose that there is such that Then the local topology covers where with taken equal to 1. Formulate two models of the labeling problem, one on call it , and one in call it It will follow that in the problem has
where and Try to argue that some sets of acyclic graphs have edge-disjoint placements in the complete graph using the reference set method.
Reconsider, briefly, the extremal example with the star and the matching. Then take the fibers as the edges of the star and the comparison orbit to be then If the fibers are the edges of the star, then is large, and so it appears possible to verify the base case. This example appears to show the necessity of the requirement on the degree sequences (that they pack in ). So, start again with a set of acyclic graphs Find a multigraph that covers a packing of the such that the degree sequence of is majorized by Use the local topology of to cover the local topology of Label such that every edge receives a label and some of the vertices receive multiple labels that correspond to source vertices of the acyclic graphs , then the multigraph is labeled and this labeling forms a probability distribution that counts the number of edge-overlaps of the when they are packed in Then the local topology of the labeled multigraph is covered by the local topology of and there is a uniform map that identifies all the fibers of this local topology.
Take equal to the expectation of duplications in the first model, and equal to the number of fibers in either model, then there is an isomorphism between the the local topology of and the local topology of The comparison orbit of is embedded isometrically in the covering of the local topology, by the local topology so that in Apply the three steps of the algorithm, and make conclusions about the labeling on See that the comparison orbit is calculated correctly in See that the comparison orbit in is one because the model is static; use the embedding of in . Do not just skip to the model in Any conclusions about the distribution of the probability at all fibers in the second model, using the reference set method and the Equality (1), apply to the distribution of edge-overlaps in because the events are between the two models before the identification is performed.
The isometric embedding of the comparison orbit permits that all the fibers can be taken identically under the product topology. In there are useful identities for using binomial expansion, and when the isomorphic embedding is applied, the various identities form the basis step of the argument. Additionally, there is the very important use of the local topology in place of the standard graph model. Establish a set of maps from the multigraph or the local topology of the multigraph to the complete graph or the local topology of the complete graph respectively. Instead of using the model of at an edge, it is possible to look at the model at each vertex.
If there is not such that then every set of maps on takes a neighborhood of into for some Then for the product topology derived to be symmetric under so that every fiber can be taken identically (so that ), it is necessary to take That is, there is that realizes a packing so that the degree sequence is majorized by a lambda-fold complete graph with lambda taken greater than 1. Use this model for the set of maps, then use all the results in the setting of with the results will not apply to the simple complete graph, and will not give necessarily edge-disjoint placements. An example will make the idea clearer. Then the following theorem applies to trees whose degree sequences pack in The approach is to label the edges and vertices of a multigraph and then use the comparison orbit in the multigraph labeling problem as it isometrically translates into a separate model. Then use the basis step from the second model, where all the calculations for beta carry isometrically. The calculations in the second model cover the first model, so that when there are no duplications across all the fibers in the final step of the induction, this statement holds in the first model. So the final step of the induction in the second model implies there are no duplications across the fibers of the second model.
The conclusions only hold for trees, because the comparison orbit in the first model is only bounded by for acyclic graphs. There are two modifications for the problem. First, use a different model of the fibers than the complete graph. That is, use the model from the local topology of the complete graph. Second, because a separate model is used apart from the usual setting, there is an allowed transformation of the variable that makes the counting different.
(Example. The Gyarfas-Lehel Conjecture) The degree sequences of the odd and even elements of pack in the degree sequences of and , respectively, where and are the maximally irregular graphs of order and This condition is strong enough to satisfy the degree condition in Step 1.
Step 1. Show that there is in the case of the example with and That is, there is with with So there is with uniform.
Step 2. Use a transmission coefficient determined by and as the event for the basis step of Step 3. First assume that the differential on the product of the two terms is continuous, and observe that gives This double counts the edge-overlap at the edge-fiber as two, and, in this case, counts the same as the 2 in the definition of (on the right-hand side). As a practical matter, note that it should actually be that the differential in the integral of is mixed, and in general, there should be some net difference on the overall solution for That is, in this case, the mixed differential counts a fixed amount that is easy to calculate, over the support because the only counting on the mixed differential occurs at the point
Then and, applying the identification, There appear to be two parts to transforming the variable: First do the transformation for Solve the definition of in for the value such that So in the first part of Step 3, suppose that the probability The next part of Step 3 is more complicated. Solve for the expectation of the random variable under the transformation, and under identification. The physical interpretation of the system is that integration occurs from one endpoint of the local topology to the other to get Therefore, to solve for under identification of the fibers, set in the upper bound of the integral on , and fix Subsequently, take the expectation. Set
The standard form of gives a factor of in front of the term, however, the fibers double cover the space, so there is an additional factor of two in the expression for Then before the identification, and as the identification changes the random variable, so that after the identification. See that the non-negativity of the variance is not used here. The transformation of the variable is such that where is the product of two identical, independent random variables, so that before the identification. Apply the identification, and rescale the value of the random variable Then in these cases, checking the degree sequence condition will bound the magnitude of The degree sequence condition will be the main constraint in applying the technique. There are cases where the reference set hypothesis does not check; in this case, the reference set hypothesis checks. In this setting, take where is the number of fibers in because the local topology of applies instead of the model in the complete graph. The cardinality of the comparison orbit is given by Count, for the value times the cardinality of the coordinates in the edge labels on Then the first coordinate is fixed by the edge incident to the oriented edge on an out-directed copy of for all except when is incident from , the source of the out-directed copy of . The choice of the first label is arbitrary with uniform distribution. The probability and the event is the event that there are no edge-overlaps in
In this case, the events are at the fibers of when they are equal to Then with uniform and with The reference set method applies with and
Step 3. The probabilities
by the rules of probability. Then there is some special consideration that allows Generally, the duplications are between fibers and have one endpoint in each of the fibers that form the edge label duplication. In this case, both of the edges in the edge label duplication are located at the same fiber of where the edge label duplication occurs. Take as the expectation, with the factor of 2 for each of the edges in the edge label duplicate pair. The factor of two covers the two edges at the same fiber of , whereas generally these two edges can be located at distinct fibers. So here, specific to this example, take for whereas normally, this equality does not apply.
Take the differential that counts the probability discretely, over the support of the desired integral in the neighborhood of 0. Then the basis step amounts to showing is larger than the integral over this support, and this holds for large
Check the reference set hypothesis using integration and Stirlingβs approximation on factorials. To count in this case, compare to the exponentiation of the integral of The integral gives Exponentiate this expression, and multiply it by to get
because This checks the reference set hypothesis, in the case of the Gyarfas-Lehel Conjecture, for sufficiently large Then the reference set method shows the Gyarfas-Lehel Conjecture holds with all generality for sufficiently large . (One of the placements of the trees has to be regarded as fixed.)
There are cases where the degree sequence condition will check, but the reference set hypothesis will not. Take, for example, the packing of spanning trees in for even In these cases, the packing of spanning trees is potentially dense. For instance, the degree sequences of three -spanning trees, each described by appending a second pendant vertex to one of the penultimate vertices in a pack in the degree sequence of . However, it is not possible to pack the three identical trees in a Pack one of the trees in the complete graph. Without loss of generality, one of the degree three vertices of the other two trees (which are identical) must pack on one of the two leaves from the pre-packed tree that are both incident the same penultimate vertex. Then the path of length three, from this vertex of degree , must have its pendant vertex packed on the same image as the degree in the original tree. Inspecting the two possible cases, which are completely determined by these conditions, yields that the leave of neither packing of the two trees is isomorphic to the third copy of the three identical trees. Now, try to apply the ideas in the example to the cases of packing spanning trees in the complete graph of even order. See that while the degree sequence condition may check, the reference set hypothesis is not going to check because
The degree-sequence packing result is not going to show packing results in general. In the case that all the sizes of the elements of the packing are identical, with a threshold of where is a linear constant that is asympotically independent of the basis step checks, but the graphs may not be acyclic and the method is not going to apply. (This usage of is independent of the other usage of In some cases, the number of fibers in an edge-disjoint placement is not going to match the number of fibers in the covering space of the local topology. Trim down to an exact cover in these cases: The generating function with the binomial distibution as a check on the basis step does not apply, as in the fashion of the example. Again, when the model applies as the target for the fibers of the problem, the probability that there is a duplication on an element of the union of the targets (fibers) covers the probability that there is a duplication in the labeling. Generally speaking, take for the results to apply to simple graphs.
3. The dynamic reference set method
Begin an attempted proof that graceful labelings exist for all classes of trees. The attempted proof is based on the reference set method, but the proof here is limited to trees. The modification of the reference set method is called the dynamic reference set method. The proof is based on total probability, instead of inclusion-exclusion. The fibers of induction are still identified in the dynamic version of the method. However, the probability space expands as the induction continues, in this version of the method, so that the calculations are different. The domain of the fibers increases, under induction, and the probability space expands. In the primary case here, the labeling is on the vertices of the graph and the random variable counts duplicated labels induced on the edges of As with the regular version of the reference set method, there must be a resolution of the graph so that the fibers resolve one by one. Meanwhile, try counting the duplications of an edge labeling at the vertices of the graph: then the vertices have to account for the duplications at both endpoints of each edge. That is, the counting for is different if the vertices are used as fibers instead of edge fibers. The results that follow here use the edges as the fibers of induction, and, as a consequence, apply only to trees and forests. Start by considering Use to express the mean of a gamma random variable in this case.
The random variable that solves the related differential equation will be a simple gamma random variable in one parameter. Usually, use for that parameter, but in this case use values from the set The overall method is analogous to the reference set method in this case. Identify the fibers in the same way, but use total probability conditions on the counting for the fibers, instead of using inclusion-exclusion. The induction in this case is similar to the standard case, and works on larger and larger numbers of fibers at each step. That is, the argument is similar, but not identical to the standard argument. Most of the proof that
is identical. Induct on the probabilities using total probability instead of the usual inclusion-exclusion argument. The argument here is still quite general, but does not follow the same rules as the argument in the setting of inclusion-exclusion. In order to be complete, review the differential equation that generates the counting functions. In this case, see that
Usually the following applies:
In the case here, there are not the added complications with expressing the solutions, because of the different formulation of the problem. Rearranging the differential equation gives that where is the probability that the number of duplicates is less than or equal to on the sum of variables. This differential equation shows that the probability density function that gives the number of free labelings is approximated closely by the gamma distribution in one variable. In this case, take
To motivate this definition, check the expression for . The counting is complicated by the requirement that for all in order to solve the differential equation. If this equation is not satisfied, find and then reset for all values This generates consistent solutions as long as for all
If that is not the case, then set larger than such that for all and set for all Use these substitutions in every case of the application of the dynamic reference set method. The induction verifies that as the domain of the vertices of expands, from one to the random variable has
as the relevant inequality. For the main example, work with a labeling problem with the edges as fibers, and domain of on the vertices.
The constant is over some resolution of That is, count the cardinality of the reference set equal to to verify the reference set hypothesis. Multiply the reciprocal of by one in order to verify the reference set hypothesis.
In some cases, the assumption that for some realization of breaks down when is not a tree. Consider a case where the labeling duplications occur on the edges, and the edges are used as the fibers of induction. Instead of adding a pendant vertex, add a vertex with two chords. The inductive orbits of the fibers differ. For example, if there is a vertex with two chords after resolving three fibers, the function for in the validation of the reference set hypothesis is
In the tree case, there are five fibers, which are all interchangeable, including the choice of the fiber in the initial stage of the induction, and all the secondary fibers have interchangeable orders of resolution. In the case where one of the fibers necessarily has chords, then there are only four inductive orbits of fibers, and one of the inductive orbits has two inductive sub-orbits. In the acyclic (tree) case the counting is clear. In the case where the graph under the labeling scheme has cycles, the counting is more convoluted, but it is clear the two cases (acyclic as opposed to cyclic) do not necessarily have identical outcomes.
When the fibers of induction are vertices, there appears to be an analogous obstruction to the inductive step in the technique, but it does not appear to be identical. For instance, in a problem where the fibers of induction are vertices, but the label duplications are on edges, the dynamic version does not appear to apply, unless the duplications at each edge are counted at each of their endpoints. This complicates the basis step of the induction.
When the fibers of induction are vertices, but the label duplications are on vertices, the dynamic version appears to apply.
For instance, in the case of Brooksβ Theorem, the basis step would normally check, but in the context of the dynamic version of the reference set method, it is not the case that So in the case of Brooksβ Theorem, the dynamic version does not appear to apply.
(Example. The Graceful Tree Conjecture) So then, validating the inductive step and in turn the reference set hypothesis will then be markedly different in these cases where we add vertices with chords instead of pendant vertices.
Step 1. The edges are identified to form the fibers of induction.
Step 2. The fibers of induction have to be edges, otherwise the count for the expectation will double for the number of endpoints of edges, in addition to the natural doubling of expectation that generally takes place. (The penalty portion of the expectation does double.) Calculate and then use to calculate the rest of the
Here, under the graceful labeling model, set and for where indexes the number of edges labeled, then set
for Fix for the sake of the argument, and see that there is a summand in the expectation for exceeding the range and that it is calculated correctly. Take
Then for all
Reset
as described. Assume that vertex labels are used only once, so there is no cycling.
Step 3. Define the random variables so that they count the number of duplicate labels on as the domain of expands. The expectations calculated are greater than or equal to the corresponding expectations for these random variables. To verify the basis case of the induction, verify that That is, show because is the cardinality of the solution space of and is the value of the probability distribution function for evaluated at 1. Use Markovβs Inequality and counting to get that
is sufficient to verify the basis case. The critical point is in the neighborhood of Then check the reference set hypothesis. Check Inequality (2) to verify the reference set hypothesis. Then the expression is roughly equivalent to
and it appears that this expression verifies the reference set hypothesis for the example is a tree of sufficiently large order.
It may be helpful to consider the cases for unions of cycles, and -graphs that are not acyclic.
To continue, first notice that there are different types of units on the random variable especially when the fibers are vertices that measure edge label duplications per vertex. In these cases, it seems to be that if there are edges per vertex, then is the probability that the duplications at independent edges, all rooted at the same vertex, are all zero. Then the edge label duplications will be zero more or less independently, depending on the problem.
4. Points of Errata
Then the problem becomes more complicated when there are multiple edges rooted at each vertex. (See some of the examples.) Then also notice that in some cases, the proportionality has larger or smaller constants of proportionality, depending on the problem. Hypothetically, suppose that if the problem measures edge label duplications, where the fibers are edges. Usually, there is a natural constant of proportionality 2, for each of the edges in a duplicate pair, and then sometimes there are different constants of proportionality in different contexts. Consider four of the problems where there are measured penalties: the Graceful Tree Conjecture, the Bollobas-Eldridge Conjecture, the Erdos-Sos Conjecture, and Sumnerβs Conjecture. In all but the Bollobas-Eldridge Conjecture, measure edge label duplicates. In the cases of Erdos-Sos and Sumnerβs Conjecture, there is a natural doubling of the penalty because the fibers are edges. In these cases, there is no factor because the -body allowance does not apply to the penalty. In the case for the Graceful Tree Conjecture, there is a natural doubling because the fibers are edges. There is no factor since this is not a -body problem. That is, in the case of the Graceful Tree Conjecture, the penalty doubles. In the case of the Bollobas-Eldridge Conjecture, the variable measures vertex duplications. There is no natural doubling (for pairs) and no natural doubling for fibers, so the constant of porportionality is 1 (with respect to the penalty for vertex-replacement). In this last case, the units of the random variable are vertex label duplications per vertex, so the dimensional analysis does not allow for any additional doubling.
In all the cases where edge label duplications are used and the fibers are vertices, there is generally a fixed map from edges to vertices that establishes the vertex-root of an edge. In these cases, the condition does not necessarily hold unless the problem is properly phrased. Usually, the factorial in the denominator of the left-hand side of the check of the reference set hypothesis is accompanied by an identical factorial in the denominator of the right-hand side of the check of the reference set hypothesis, and these identical terms cancel, so that there is an abbreviated check of the reference set hypothesis. In order to do this, the orbits of the fibers are required to accumulate multiplicatively to form the factorial factor.
Try packing graphs or, equivalently, consider a labeling problem where the duplicates are between graphs instead of among graphs. The cardinality of the comparison orbit in is the total number of categories for fiber duplication, in these cases. That is, the comparison orbit is the number of categories which are considered for counting duplicate pairs. (In some cases, this is one-sided because the relation of a comparison orbit is not transitive: If and it is not necessarily the case that It is the case that the cardinalities of the comparison orbits do not have to be identical either.)
Resolve two cycles as in the case of -labelings. Then the factorial count for the order of the permutations of the fibers is not necessarily . The concern may be that the number of inductive orbits on the resolution of the graph is path-dependent, and that no resolution gives The probability is established by accounting for the number of edges mapped to the fiber under the map, as long as the fibers are vertices. If the fibers are edges, then the inductive orbits at a fiber of the resolution may fall into inductive sub-orbits, and this can cause some of the hypotheses of the reference set method to fail. The maximum factorial count for two disjoint cycles should be In this case, do not just apply this value to the denominator of the term in the reference set hypothesis check. Because the value is not exactly the reference set method does not apply at all. For this reason, it appears that the -labeling and -labeling results apply only to trees and forests (for all orders and sizes identically). It appears that in vertex labeling problems, when duplicates are counted on edges, and the edges are used as the fibers, the comparison orbit is not well-defined for cyclic graphs. That is, it appears that in these problems, the comparison orbits are only well-defined when the elements of the labeling problem are limited to acyclic graphs.
That is, there is an obstruction for vertex labeling problems, with vertices as fibers, where is taken over edges. In this case, the fibers can induce duplication units so that the ratio of fibers to edges cannot necessarily be taken in the case of unicyclic graphs. These restrictions do not apply to the Vizing Theorem or the Total Labeling Conjecture. In neither of these two cases do the labels on vertices induce labels on edges, though in the one case the labels on vertices may affect the comparison orbit on the edges.
The Gyarfas-Lehel result does not appear to apply to graphs generally, because, even though there is a separate model for the results, and the labeling problem is different, the comparison orbit is only defined as described for acyclic graphs, the label on the endpoint of each edge has possible values.
Then the argument on the Tree Packing Conjecture has several complications that are not present in many of the other examples. Because there is not a clear example of a constructive technique that verifies the conclusions here regarding the Tree Packing Conjecture, these results should be viewed very critically. The proof included here on the Gyarfas-Lehel Conjecture assumes the local topology without fixing both the endpoints of the fibers. This technique of using the local topology is different from the other examples. Using a uniform map and a covering from the local topology of the complete graph, the events in the two distributions can be mapped The conclusions given about the conjecture should follow because there is an isomorphism between the two local topologies, and because the cardinality of the comparison orbit is carried isometrically from model to model. Apply the conclusion on the one model in the other model. The second model would normally have a static comparison orbit and the basis step is easier to do in the second model. Analogously, the cardinality of the comparison orbit is in the case of Bollobas-Eldridge, Erdos-Sos and Sumnerβs Conjecture if vertex-replacement is not allowed. However, in these three cases replacement is allowed and so the comparison orbit is Then the Erdos-Sos and Sumner Conjecture results apply to trees, but not necessarily other graphs. (In the case of Erdos-Sos, check the results that hold in the case ; the analysis suggests that there are some results.) The set-up for Gyarfas-Lehel gives a closed form for for a model of the problem labeling the edges of a multigraph (and some vertices) where is some packing of the graphs in a multigraph whose degree sequence is majorized by
In most cases, the reference set method gives a set of three initial conditions that are sufficient to demonstrate that the amount exceeds the threshold
If, in addition, this threshold is greater than the reciprocal of the cardinality of the reference set, then there is an instance of that realizes the probability that the labeling problem has an instance with no duplications. (Address the inherent problem with describing probabilities that are greater than zero, which have no elements in the sample set that realize those probabilities.) This appears contrary to some of the previous developments, particularly in random graph theory. The problem may be partly semantic. Because the induction begins by modeling the original distribution with a function relative to a continuous differential, there can be fractional probabilities in the derived distributions that are not realized by the original probability distribution. When the probability exceeds certain thresholds, that is not the case. Only probabilities that have integer numerators divided by are properly realized in these problems with respect to the bounds on the probability distributions derived here. Call this last condition proper realization. In order to show the basis step of the third step of the reference set method is sufficient, use that
on any support where
with respect to any mixed counting differential that has
when is any non-negative integer. Consider a sequence of functions that converges point-wise to with discrete masses at Using weight the discrete points 1 at each discrete point in and on every point of the support
Take the limit of the sequence of functions as . Then the point-wise limit of the sequence is infinitely differentiable on the full support in the following sense: The limit of the sequence of functions is a cumulative density function under normalization, and converges point-wise to . Furthermore, the functions can be approximated by dirac-delta functions to get the differentiability of the cumulative distributions in the sense that (under proper realization) so that for all and upon induction,
Take as the normalizer of . That is, let be taken equal to the integral of over the full support with respect to Then if
it follows that, for instance,
as long as all the conditions from Step 1 and Step 2 of the reference set method are satisfied. The idea is that on some small support