### v19

1. Introduction
The ideas here are meant to give a general procedure for calculating an optimal sufficient value for the number of labels required in a standard labeling problem. The techniques apply to vertex labeling and edge labeling problems. Usually, the calculation will give a rough estimate of the optimal sufficient value, rather than an exact value. However, the calculation should bound the optimal value from above in most applications. There are several steps to proving the main result, which is that given three sufficient conditions, it is possible to calculate upper bounds on the number of labels required in the general labeling problem. The first sufficient condition is that there is a uniform map that carries the fibers to one another transitively. Then, in addition, there is a basis step to an inductive argument. Usually, this step is verified by applying Markovβs Inequality to an expectation that is calculated for each of the individual fibers, before performing the identification allowed by the existence of the uniform map. Once the basis step is verified, apply induction. The induction implies that there is a simple test for the verification of a fixed upper bound on the number of required labels. The required step here is verification of the reference set hypothesis; show that the reciprocal of the sample set for the number of labelings is less than the probability that there is a labeling without collisions. The probability, based on an expectation parameterized by the projected required strength of the labeling, is calculated using integration given by the induction from the basis step of the induction. There is a correlation between the size of an expected value, and the existence of a labeling without collisions, in the framework of each individual labeling problem.
Look more closely at packings with exactly two elements, which are called -body problems. For most -body problems, depending on the phrasing of the problem, formulate results without using the required hypothesis that there is a placement of the two graphs so that the sum of the degree sequences is majorized by the degree sequence of the complete graph. There are other hypotheses used as substitutes for the degree packing hypothesis. In these cases, there is a -body allowance that permits the use of without the natural doubling of expectation that is normally required in figuring
(Example. The Bollobas-Eldridge Conjecture) The Bollobas-Eldridge Conjecture is formally stated that if then and have an edge-disjoint placement in . There is a very strong analogue. There are complications in every step of the method with regards to analyzing the proof of the result. All the complications and some of the nuances of the method in this setting are explained next. The maps have to allow label replacement to get the result. So, there is a penalty for vertex duplication in the expectation. Step 3 is complicated by the assessment of the penalty for vertex label duplication and in the check of the reference set hypothesis. Additionally, the calculation for the value of is somewhat complicated.
There are some obvious extremal examples such as Furthermore, the examples of disjoint collections of suitably ordered complete graphs are extremal in terms of the problem statement. There is a previous outline of an approximation of the result with previous ideas: the reference set method shows an approximation of the Sauer-Spencer criteria. The previous result can be improved to show a closer approximation of the desired conjecture in most cases. Outline the problem as a vertex labeling problem (as opposed to an edge labeling problem where the edge label duplications are counted at the vertex fibers) and form where the terms are the expected number of edge label duplications between and in the problem. Additionally, in this case, is set
Step 1. Use a map from to the looped complete graph where and a map from to the same target. The symmetric map from into is uniform, as is the map from The map allows replacement; the application requires that the maps be injections, and therefore, there is a penalty assessed on the random variable and the expectation of the random variable. Checking the degree sequence requirement is not required in this case, because there are only two graphs in the packing.
Step 2. There are two parts to the stage in the problem. There are two parts to . The first part comes from the expectation in the problem. The second part is a penalty for the duplication of vertex labels. Consider all the placements of and in where replacement is allowed on the vertex labels of and (In this way the problem is highly atypical because loops are allowed and then counted in the random variable.) In this case is the expected number of edge label intersections in closed neighborhoods from against closed neighborhoods of the vertices taken over the range of and Count the number of ways edges by dividing into three parts and considering all the separate cases from various sets of closed vertex neighborhoods: Consider first the vertices are mapped to the same target in and Consider and Take This should be clear from simple counting. Use in the denominator of the multiplier instead of because the labeling allows replacement, and so the product is uniform over choices instead of . The degree of the vertices is instead of as in a standard complete graph. There is a penalty for vertex label duplication, but it is not applied in this part of the expectation. Consider second the total expectation that both and Compare the closed neighborhood of and This expectation is The counting here is more complicated, but most of the main ideas can be gathered from the expression for . Consider third the vertices are mapped to the same target in and Consider and The third component of the expectation is The counts appear correct.
Set Sum over all ordered pairs of vertices from and respectively, and each vertex labeling is counted twice, once for each endpoint of the edge. Notice the allowance for the -body problem. Then there is a penalty for allowing vertex-replacement: Take the penalty equal to Do not double the penalty for the two elements of every duplication, and there is no natural doubling because the fibers are vertices. The -body allowance does not apply to the penalty. The number of fibers and the comparison orbit of the fibers in the penalty have identical cardinality. Take Count the vertices of as fibers, and each fiber has the comparison orbit cardinality (Do not count the vertex fibers as ) These are most of the difficult steps.
Step 3. In the reformulation of the problem, this step holds with and Take and use Markovβs Inequality. See that if then the basis step is satisfied. To check the reference set hypothesis, use . Count the number of labelings with replacement of both graphs. Then the value is about and is Try to use the edges as the fibers of induction, and get a function that is hyper-exponential in some base, whereas is only when replacement is allowed.
Continue by addressing two of the labeling problems that involve finding resolutions of trees in graphs and digraphs. The applications that follow appear mostly limited to trees.
(Example. Erdos-Sos) Here, the -body specification applies to the first part of the expectation. Both elements of the duplications on the vertex labels are internal to the first graph, so there is no -body allowance for the penalty portion of the expectation.
Step 1. The fibers of the problem are the edges. The labeling is on the vertices. The problem requires that the graph under resolution is a tree or forest. Because there are only two graphs in the packing, it is not necessary to verify the degree sequence condition. Use the -body allowance, and take the edges as the fibers of induction. It appears necessary to take a penalty in the expectation for vertex label duplication, and use
Step 2. Here, let the tree have order . Assume the host graph has average degree taken slightly larger than the requirement of the problem. Then Then for this component is and For the penalty portion, Then for this portion is applied as and the -body short-cut does not apply: The penalty is The penalty portion of the expectation doubles. So then
Step 3. There is a sufficiently large average degree in the host graph where the basis step will check. See that it is in the neighborhood of the specifications of the problem. Estimate the value of as using the expression for The idea is to show that Erdos-Sos is asymptotically accurate. The check of the reference set hypothesis is because in most cases is sufficiently smaller than . If beta is not sufficiently small, adjust the density of the graph slightly. Do not choose much smaller than because it is not required in the basis step of induction. Then the basis step and the reference set hypothesis both check, and these calculations give a rough form of the Erdos-Sos Conjecture.
(Example. Sumnerβs Conjecture) Here, the -body specification applies to the first part of the expectation, but not in the second part, because all the collisions on labels of the vertex set are internal to the problem.
Step 1. Skip the details for this step. The problem is phrased as a packing problem.
Step 2. Try to pack in the tournament Take Take Then Take, for the penalty Take The -body allowance does not apply to and so The penalty portion does not naturally double, but because the fibers have two endpoints, the penalty portion doubles.
Step 3. The basis step does not check unless the size of the tree is reset to and is sufficiently smaller than In these cases, the basis step checks because To check the reference set hypothesis, use the reference set because replacement is allowed.
For the -body problems, use the edges of the smaller graph as the fibers of induction, and usually maintain that the smaller graph in the placement is acyclic. In all these cases, it is usually necessary to take the penalty for vertex label duplication, and when the edges are the fibers of induction, the penalty doubles. To apply Markovβs Inequality to the basis case, compare some form of given by the individual fibers to the value The calculation for will generally vary from problem to problem, and is not always easy to establish. Try to extend the Erdos-Sos result to a -factor, using the edges as fibers. Then the reference set hypothesis will not check unless The necessary restriction on in this example excludes any new results, and gives standard results within accepted parameters. Similarly, use the vertices as fibers and get results for cycles. Show gives and gives Try to extend the Sumner Conjecture to cycles. The phrasing of the problem as a vertex fiber problem will not permit a check of the reference set hypothesis.
(Example. Turanβs Theorem) Try deriving an approximation to Turanβs Theorem: The graph if and The closest approximation here is that packs in when the complement of has with In this case, do not apply the -body allowance; do not take the full penalty for vertex label duplication.
Step 1. This step is carried out using the symmetric map from the vertex set of to This map is uniform on the edges of the selected. Because the selected graph and the target graph are isomorphic to , it is clear the map will be uniform. This example is unusual because Do not check the degree sequence requirement in this example. Use the vertices as the fibers.
Step 2. The calculation of does not use the -body allowance. The constant of proportionality is 4 in this case. The duplication units are edges and the fibers are vertices. Use as the density of null edges in the graph that the vertices are chosen from. Use the formula to check the base case. (Pick about at a vertex of if is a small fraction close to and using exponentiation. This calculation appears to hold in general for cliques, but is not applied in this setting.) Take Then and and the value when See the case where the density of edges in the maximal complementary graph should be on the order of -factor: the choice of here as gives that is about
Step 3. Verify the basis step. Take the fibers of induction to be the vertices of Then take Again, choose and sufficiently small. Then is the calculation for the reference set hypothesis. By choice of sufficiently small choice is taken sufficiently smaller than
where and The calculation measures that the probability of a labeled graph in vertex labels intersected with and having no empty edges as greater than the reciprocal of the number of vertex labelings of with labels (where only vertices of are labeled). This would verify an approximate of Turanβs Theorem by the reference set hypothesis for suitably small The point is to drive smaller than because it is necessary to check the reference set hypothesis. Here and larger choices of give larger choices of The resultant larger choices of are more likely to disobey the reference set hypothesis. This makes intuitive sense because more null edges would decrease the likelihood of giving a choice of a complete The value is approximately the largest value that checks in the basis step. The value is in the denominator of the value for and is taken It has to be that is satisfied. The formulation of the theorem does not give information about the unique extremal graphs generated by the problem. Apply Markovβs Inequality in this context. Here, the choice of is limited by the basis case and the value of does not appear optimal in this problem.
Summarize some of the techniques that are available for edge-disjoint placement problems. First, to pack for generally check the degree sequence condition, use a covering space with a multigraph, and possibly use a transformation on the random variable. (In this case, the graphs have to be acyclic.) Then, to pack and there are several different formulations of the problem, most of which rely on the -body allowance in some way, and some of which do not require the degree sequence condition as an initial hypothesis. First, take the edges as the fibers of induction. In this case, take the penalty for vertex label duplication and use the cardinality of the comparison orbit Do not check the degree sequence requirement. To apply this case, the graph has to be acyclic. Next, use the idea from the example on the Bollobas-Eldridge Conjecture.
In this case, it does not appear necessary to check the degree sequence requirement explicity, but add to the expectation to account for the vertex label duplication because of the phrasing of the problem. Third, use the vertices as fibers, with the added penalty for vertex label duplication, with similar conditions to the edge fiber case, so that neither graph has to be acyclic, but formulate in terms of and take Additionally, in this case, the penalty for vertex duplication is based on instead of the order of the graph Here, it is not necessary to check the degree sequence condition separately. Fourth, take the vertices of one graph as the fibers of induction, and skip checking the degree sequence requirement. In this case, set the cardinality of the comparison orbit equal to the product of the cardinalities of the vertex sets of the individual graphs:
This formulation does not appear to have any optimal applications. This last formulation also appears to apply in the case but again, does not seem to have any optimal applications.
2. Degree sequences and edge-disjoint placements
First, notice the following: Suppose the inequality is the only inequality available in proving the base case of Step 3. Then it may be that the smallest inequality that makes sense is actually for The strongest conclusion on using this basis step on the reference set method is
This is important, because it generally means problems cannot be solved with the sample space out of factorials by extending the support from to
There are additional complications when the problem is an edge-disjoint packing problem. There are alternative phrasings for edge-disjoint packing problems: One alternative is to use the fibers as edges and to use the entire range of the complete graph as the comparison orbit. Then and let where and The expectation
so that in the case that the packing is dense. Try the case of two spanning stars. Since the degree sequences do not pack, it is not possible to pack the two stars in an edge-disjoint fashion; try to verify that there is a packing in some of the other situations without checking the degree sequence condition. Use the full range of the edge set of the complete graph as the comparison orbit, or formulate the problem alternatively. One alternative is to use the vertices as fibers and to use the cardinality of the cross product of the vertex sets as the comparison orbit. Generally, there is the identical difficulty when the number of elements in the packing is in the case of two spanning stars, show that this parameterization will not give false positive results. Then and where and where the packing has elements. In the case of two stars, take and Then and clearly the method is impractical. There are similar difficulties when the problem is rephrased for different elements in the graph-packing. Both alternatives for the formulation of the random variable present obstacles that are realized to some extent by extremal examples. Consider the case of the star and the matching: In that case, given this parameterization, so that when is the order of the star. There are various ways of considering the same problem for packing sets of graphs. One of the goals is to set up the problem so that the ideas can be applied in a wide range of settings. The following idea for setting up the problem is more flexible, but has some added requirements that are difficult to satisfy.
The graph packing problem can be resolved in some cases by adding a degree sequence packing requirement. Proceed by taking and then labeling the resultant packing as a multigraph that is labeled and re-labeled. Let be some packing of a set of acyclic graphs Generally work in instead of to avoid problems. Take as usual. Suppose that there is such that Then the local topology covers where with taken equal to 1. Formulate two models of the labeling problem, one on call it , and one in call it It will follow that in the problem has
where and Try to argue that some sets of acyclic graphs have edge-disjoint placements in the complete graph using the reference set method.
Reconsider, briefly, the extremal example with the star and the matching. Then take the fibers as the edges of the star and the comparison orbit to be then If the fibers are the edges of the star, then is large, and so it appears possible to verify the base case. This example appears to show the necessity of the requirement on the degree sequences (that they pack in ). So, start again with a set of acyclic graphs Find a multigraph that covers a packing of the such that the degree sequence of is majorized by Use the local topology of to cover the local topology of Label such that every edge receives a label and some of the vertices receive multiple labels that correspond to source vertices of the acyclic graphs , then the multigraph is labeled and this labeling forms a probability distribution that counts the number of edge-overlaps of the when they are packed in Then the local topology of the labeled multigraph is covered by the local topology of and there is a uniform map that identifies all the fibers of this local topology.
Take equal to the expectation of duplications in the first model, and equal to the number of fibers in either model, then there is an isomorphism between the the local topology of and the local topology of The comparison orbit of is embedded isometrically in the covering of the local topology, by the local topology so that in Apply the three steps of the algorithm, and make conclusions about the labeling on See that the comparison orbit is calculated correctly in See that the comparison orbit in is one because the model is static; use the embedding of in . Do not just skip to the model in Any conclusions about the distribution of the probability at all fibers in the second model, using the reference set method and the Equality (1), apply to the distribution of edge-overlaps in because the events are between the two models before the identification is performed.
The isometric embedding of the comparison orbit permits that the cardinality of the comparison orbit is well-defined and is calculated correctly. In there are useful identities for using binomial expansion, and when the isomorphic embedding is applied, the various identities form the basis step of the argument. Additionally, there is the very important use of the local topology in place of the standard graph model. Establish a set of maps from the multigraph or the local topology of the multigraph to the complete graph or the local topology of the complete graph respectively. Instead of using the model of at an edge, it is possible to look at the model at each vertex.
If there is not such that then every set of maps on takes a neighborhood of into for some That is, in some instances, every multigraph that realizes a packing is majorized by a lambda-fold complete graph with lambda taken greater than 1. Try using this model for the set of maps, so that all the results in the setting of have the results will not apply to the simple complete graph, and will not give necessarily edge-disjoint placements. If there is such that with then there is a valid covering space using the local topology of the complete graph. In this case, invoke induction using the cardinality of the comparison orbit, so that the reference set hypothesis applies. An example will make the idea clearer. The following theorem applies to trees whose degree sequences pack in The approach is to label the edges and vertices of a multigraph and then use the comparison orbit in the multigraph labeling problem as it isometrically translates into a separate model. Then use the basis step from the second model, where all the calculations for beta carry isometrically. The calculations in the second model cover the first model, so that when there are no duplications across all the fibers in the final step of the induction, this statement holds in the first model. That is, the final step of the induction in the second model implies there are no duplications across the fibers of the first model. The conclusions only hold for trees, because the comparison orbit in the first model is only bounded by for acyclic graphs. There are two modifications for the problem.
First, use a different model of the fibers than the complete graph. That is, use the model from the local topology of the complete graph. Second, because a separate model is used apart from the usual setting, there is an allowed transformation of the variable that makes the counting different.
(Example. The Gyarfas-Lehel Conjecture) The degree sequences of the odd and even elements of pack in the degree sequences of and , respectively, where and are the maximally irregular graphs of order and This condition is strong enough to satisfy the degree condition in Step 1.
Step 1. Show that there is in the case of the example with and That is, there is with with So there is with uniform.
Step 2. Use a transmission coefficient determined by and as the event for the basis step of Step 3. First assume that the differential on the product of the two terms is continuous, and observe that gives This double counts the edge-overlap at the edge-fiber as two, and, in this case, counts the same as the 2 in the definition of (on the right-hand side). As a practical matter, note that it should actually be that the differential in the integral of is mixed, and in general, there should be some net difference on the overall solution for That is, in this case, the mixed differential counts a fixed amount that is easy to calculate, over the support because the only counting on the mixed differential occurs at the point
Then and, applying the identification, There appear to be two parts to transforming the variable: First do the transformation for Solve the definition of in for the value such that So in the first part of Step 3, suppose that the probability The next part of Step 3 is more complicated. Solve for the expectation of the random variable under the transformation, and under identification. The physical interpretation of the system is that integration occurs from one endpoint of the local topology to the other to get Therefore, to solve for under identification of the fibers, set in the upper bound of the integral on , and fix Subsequently, take the expectation. Set
The standard form of gives a factor of in front of the term, however, the fibers double cover the space, so there is an additional factor of two in the expression for Then before the identification, and as the identification changes the random variable, so that after the identification. See that the non-negativity of the variance is used here. The transformation of the variable is such that where is the product of two copies of an identical random variable, so that before the identification. Apply the identification, and rescale the value of the random variable Then in these cases, checking the degree sequence condition will bound the magnitude of The degree sequence condition will be the main constraint in applying the technique. There are cases where the reference set hypothesis does not check; in this case, the reference set hypothesis checks. In this setting, take where is the number of fibers in because the local topology of applies instead of the model in the complete graph. The cardinality of the comparison orbit is given by Count, for the value times the cardinality of the coordinates in the edge labels on Then the first coordinate is fixed by the edge incident to the oriented edge on an out-directed copy of for all except when is incident from , the source of the out-directed copy of . The choice of the first label is arbitrary with uniform distribution. The probability and the event is the event that there are no edge-overlaps in
In this case, the events are at the fibers of when they are equal to Then with uniform and with The reference set method applies with and
Step 3. The probabilities
by the rules of probability. Then there is some special consideration that allows Generally, the duplications are between fibers and have one endpoint in each of the fibers that form the edge label duplication. In this case, both of the edges in the edge label duplication are located at the same fiber of where the edge label duplication occurs. Take as the expectation, with the factor of 2 for each of the edges in the edge label duplicate pair. The factor of two covers the two edges at the same fiber of , whereas generally these two edges can be located at distinct fibers. So here, specific to this example, take for whereas normally, this equality does not apply.
Take the differential that counts the probability discretely, over the support of the desired integral in the neighborhood of 0. Then the basis step amounts to showing is larger than the integral over this support, and this holds for large
Check the reference set hypothesis using integration and Stirlingβs approximation on factorials. To count in this case, compare to the exponentiation of the integral of The integral gives Exponentiate this expression, and multiply it by to get
because This checks the reference set hypothesis, in the case of the Gyarfas-Lehel Conjecture, for sufficiently large Then the reference set method shows the Gyarfas-Lehel Conjecture holds with all generality for sufficiently large . (One of the placements of the trees has to be regarded as fixed.)
There are cases where the degree sequence condition will check, but the reference set hypothesis will not. Take, for example, the packing of spanning trees in for even In these cases, the packing of spanning trees is potentially dense. For instance, the degree sequences of three -spanning trees, each described by appending a second pendant vertex to one of the penultimate vertices in a pack in the degree sequence of . However, it is not possible to pack the three identical trees in a Pack one of the trees in the complete graph. Without loss of generality, one of the degree three vertices of the other two trees (which are identical) must pack on one of the two leaves from the pre-packed tree that are both incident the same penultimate vertex. Then the path of length three, from this vertex of degree , must have its pendant vertex packed on the same image as the degree in the original tree. Inspecting the two possible cases, which are completely determined by these conditions, yields that the leave of neither packing of the two trees is isomorphic to the third copy of the three identical trees. Now, try to apply the ideas in the example to the cases of packing spanning trees in the complete graph of even order. See that while the degree sequence condition may check, the reference set hypothesis is not going to check because
The degree-sequence packing result is not going to show packing results in general. In the case that all the sizes of the elements of the packing are identical, with a threshold of where is a linear constant that is asympotically independent of the basis step checks, but the graphs may not be acyclic and the method is not going to apply. In some cases, the number of fibers in an edge-disjoint placement is not going to match the number of fibers in the covering space of the local topology. Trim down to an exact cover in these cases: The generating function with the binomial distibution as a check on the basis step does not apply, as in the fashion of the example. Again, when the model applies as the target for the fibers of the problem, the probability that there is a duplication on an element of the union of the targets (fibers) covers the probability that there is a duplication in the labeling. Generally speaking, take for the results to apply to simple graphs.
3. The dynamic reference set method
The ideas behind the reference set method can be used to formulate a proof technique that is more flexible in some contexts than the standard method. The modification of the reference set method is called the dynamic reference set method. Set
When figuring the expectation from do not divide by the number of fibers, or multiply by the cardinality of the comparison orbit. Simply fix From this starting point, the argument is similar, but not identical to the standard argument. In order to be complete, review the differential equation that generates the counting functions. In this case, see that
where Usually the following applies (with normally taken as ):
The three steps still apply, but Step 1 is more important, because the dynamic reference method sometimes permits the application of to satisfy the basis step. The contrast in the two forms of the primary differential equation is reflected in the types of applications.
The idea behind the dynamic reference set method is that if a system has a total expectation of then the system can be modeled by a set of random variables with In the dynamic reference set method, the base case can have a non-trivial solution space, and this consideration changes the counting. That is, the dynamic reference set method sometimes permits a simpler check for the basis case, and the case when has many helpful applications. In addition to the fact that the lower bound for is sometimes smaller in the dynamic version, the basis case only requires that one fiber satisfy the basis step. The standard method requires that all fibers satisfy the basis step. The main differential equation uses total probability, so the fibers are not identified in exactly the same way. Check that
in order to verify the reference set hypothesis. Particularly in the case considered here, it is required that the comparison orbit is well-defined, which will require in some cases that the graph under resolution is acyclic. For most of the problems where the fibers are vertices, the hypothesis on the comparison orbit is satisfied.
For instance, in the case of Brooksβ Theorem, the basis step of the dynamic reference set method does not check unless The inequality has to use the value taken On the other hand, when the value is sometimes too large to derive results for Brooksβ Theorem. Consider what happens when the value for approaches zero, and approaches some value equal to
(Example. The Graceful Tree Conjecture) Try to apply the dynamic reference set method to the popular conjecture about graceful labelings. The application here is limited to graphs that are acyclic.
Step 1. The edges are identified to form the fibers of induction.
Step 2. The fibers of induction have to be edges, otherwise the count for the expectation will double for the number of endpoints of edges, in addition to the natural doubling of expectation that generally takes place. (The penalty portion of the expectation does double.) Calculate and then use to calculate the rest of the Here, under the graceful labeling model, the total penalty for vertex duplication is Set Measure
Fix for the sake of the argument. Use the expectation of the edge degree in the test for the basis step. The application of the expectation is permitted in the dynamic reference set method, because the labeling scheme is uniform to the labels, and the basis step has to hold for at least one of the fibers in the graph. Vary the value of in the interval The maximum value for is attained at the left end-point of the interval. When the induction is performed, set The number of fibers is set equal to The value for
Step 3. Evaluate Check the basis step first. That is, verify that
for sufficiently large. The inequality is closest when is maximized. Since the labels are uniform, take the two probabilities independently to get That is, the calculation for is independent of the labels chosen for the vertices that form the end-points of an arbitrary edge. The value can be taken The test for the reference set hypothesis is weaker than
which holds for sufficiently large, since and The value is sufficient to check the basis case, and the reference set hypothesis checks under this condition, as well.
Notice that, in general, the test for the basis step is
It may be helpful to consider the cases for unions of cycles, and -graphs that are not acyclic. To continue, first notice that there are different types of units on the random variable especially when the fibers are vertices that measure edge label duplications per vertex. In these cases, it seems to be that if there are edges per vertex, then is the probability that the duplications at independent edges, all rooted at the same vertex, are all zero. Then the edge label duplications will be zero more or less independently, depending on the problem.
4. Points of Errata
Then the problem becomes more complicated when there are multiple edges rooted at each vertex. (See some of the examples.) Then also notice that in some cases, the proportionality has larger or smaller constants of proportionality, depending on the problem. Hypothetically, suppose that if the problem measures edge label duplications, where the fibers are edges. Usually, there is a natural constant of proportionality 2, for each of the edges in a duplicate pair, and then sometimes there are different constants of proportionality in different contexts. Consider four of the problems where there are measured penalties: the Graceful Tree Conjecture, the Bollobas-Eldridge Conjecture, the Erdos-Sos Conjecture, and Sumnerβs Conjecture. In all but the Bollobas-Eldridge Conjecture, measure edge label duplicates. In the cases of Erdos-Sos and Sumnerβs Conjecture, there is a natural doubling of the penalty because the fibers are edges. In these cases, there is no factor because the -body allowance does not apply to the penalty. In the case for the Graceful Tree Conjecture, there is a natural doubling because the fibers are edges. There is no factor since this is not a -body problem. That is, in the case of the Graceful Tree Conjecture, the penalty doubles. In the case of the Bollobas-Eldridge Conjecture, the variable measures vertex duplications. There is no natural doubling (for pairs) and no natural doubling for fibers, so the constant of porportionality is 1 (with respect to the penalty for vertex-replacement). In this last case, the units of the random variable are vertex label duplications per vertex, so the dimensional analysis does not allow for any additional doubling.
In some particular labeling problems, the comparison orbits are only well-defined when the elements of the labeling problem are limited to acyclic graphs. The Gyarfas-Lehel result does not appear to apply to graphs generally, because, even though there is a separate model for the results, and the labeling problem is different, the comparison orbit is only defined as described for acyclic graphs, the label on the endpoint of each edge has possible values. Then the argument on the Tree Packing Conjecture has several complications that are not present in many of the other examples. Because there is not a clear example of a constructive technique that verifies the conclusions here regarding the Tree Packing Conjecture, these results should be viewed very critically. The proof included here on the Gyarfas-Lehel Conjecture assumes the local topology without fixing both the endpoints of the fibers. This technique of using the local topology is different from the other examples. Using a uniform map and a covering from the local topology of the complete graph, the events in the two distributions can be mapped The conclusions given about the conjecture should follow because there is an isomorphism between the two local topologies, and because the cardinality of the comparison orbit is carried isometrically from model to model. Apply the conclusion on the one model in the other model. The standard model would normally have a static comparison orbit and the basis step is easier to do in the standard model. Analogously, the cardinality of the comparison orbit is in the case of Bollobas-Eldridge, Erdos-Sos and Sumnerβs Conjecture if vertex-replacement is not allowed. However, in these three cases replacement is allowed and so the comparison orbit is Then the Erdos-Sos and Sumner Conjecture results apply to trees, but not necessarily other graphs. (In the case of Erdos-Sos, check the results that hold in the case ; the analysis suggests that there are some results.) The set-up for Gyarfas-Lehel gives a closed form for for a model of the problem labeling the edges of a multigraph (and some vertices) where is some packing of the graphs in a multigraph whose degree sequence is majorized by The reference set method gives a set of initial conditions that are sufficient to show that the amount exceeds the threshold
If, in addition, this threshold is greater than the reciprocal of the cardinality of the reference set, then there is an instance of that realizes the probability that the labeling problem has an instance with no duplications. (Address the inherent problem with describing probabilities that are greater than zero, which have no elements in the sample set that realize those probabilities.) This appears contrary to some of the previous developments, particularly in random graph theory. The problem may be partly semantic. Because the induction begins by modeling the original distribution with a function relative to a continuous differential, there can be fractional probabilities in the derived distributions that are not realized by the original probability distribution. When the probability exceeds certain thresholds, that is not the case. Only probabilities that have integer numerators divided by are properly realized in these problems with respect to the bounds on the probability distributions derived here. Call this last condition proper realization. In order to show the basis step of the third step of the reference set method is sufficient, use that
on any support where
with respect to any mixed counting differential that has
when is any non-negative integer. Consider a sequence of functions which are everywhere zero on except point masses at and the point mass where Weight the points 1, at each of the point masses, with on and zero elsewhere on the support. On the function Take the limit of the sequence of functions as . The point-wise limit of the sequence of functions is infinitely differentiable, in the sense that it is a cumulative density function under normalization, and the limit under is
That is, the functions can be replaced by dirac-delta functions to get the differentiability of a cumulative distribution in the sense that (under proper realization) so that for all and upon induction,
The distribution on for is not explicitly defined. The distribution on the will vary, and the only requirement that needs to be satisfied is that It follows that from the original differential equation on the Next, It follows that
as long as all the conditions from Step 1 and Step 2 of the reference set method are satisfied, and The idea is that on some small support
There appear to be several errors in [4]. For instance, in [4], it is suggested that
is sufficient to verify the reference set hypothesis in the dynamic case, and that whenever the comparison orbit is well-defined for the labeling problem, This ignores the possibility that and does not appear to be a useful way to think about the theory, because, for instance, in some cases and This last intended application still lends itself to interpretations where is some product of the cardinality of inductive orbits. That is, the product in the denominator of the right hand side of Inequality (2) is a -falling expression in some large value, where because the solution space of the base set is taken to be non-trivial in the dynamic reference set method. Then the ratio of on the left hand side of the proposed Inequality (3) is meant to be the ratio of the cardinality of the solution space of the base set with the product of cardinalities of the inductive orbits in [4]. However, see that the cardinality of the solution space of the base set is trivial in some cases, or much larger or smaller than so that this interpretation of the theory does not appear consistent. Rather, it appears that the Inequality (3) does not apply, but that the dynamic reference set method requires that the comparison orbit is well-defined, and that the product of the cardinality of the inductive orbits of a graph under resolution is
One of the advantages of the dynamic reference set method is that the base case only requires that one fiber pass the test for the basis case, as long as the labeling is uniform to the labels. That is, only one fiber has to pass the test for the basis case, as long as the explicit choice of labels applied to a fiber does not affect the probability that the label on the particular fiber has a duplication. Instead of using inclusion-exclusion, the dynamic reference set method uses total probability. Therefore, once the basis step is demonstrated for one fiber, the other fibers can be added sequentially. In order, subtract from the probability that the duplications over some set of fibers that contains the fibers in the induction has a label duplication, conditioned to the probability that the previously collected fibers do not register a duplication: Subtract The result is that the dynamic reference set method gives some resolution of a graph that has no label duplications.